Eddie's age is 4 times Marco's. In 10 years Eddie will be twice as old as Marco will be. How old is each now?
Thank you.

Answer by richwmiller(17219)

e=4m
e+10=2*(m+10)
4m+10=2m+20
2m=10
m=5
e=20
check
30=2*15
ok